# The Fragile Nature of Road Transportation Systems

### 01 Feb 2024 v1

@@ Abstract
Major cities worldwide experience problems with the performance of their road transportation systems. The continuous increase in traffic demand presents a substantial challenge to the optimal operation of urban road networks and the efficiency of traffic control strategies. Although robust and resilient transportation systems have been extensively researched over the past decades, their performance under an ever-growing traffic demand can still be questionable. The operation of transportation systems is widely believed to display fragile property, i.e., the loss in performance increases exponentially with the linearly increasing magnitude of disruptions, which undermines their continuous operation. The risk engineering community is now embracing the novel concept of (anti-)fragility, which enables systems to learn from historical disruptions and exhibit improved performance as disruption levels reach unprecedented magnitudes. In this study, we demonstrate the fragile nature of road transportation systems when faced with either demand or supply disruptions. First, we conducted a rigorous mathematical analysis to theoretically establish the fragile nature of the systems. Subsequently, by taking into account real-world stochasticity, we implemented a numerical simulation with realistic network data to bridge the gap between the theoretical proof and the real-world operations, to study the impact of uncertainty on the fragile property of the systems. This work aims to help researchers better comprehend the necessity to explicitly consider antifragile design toward the application of future traffic control strategies, coping with constantly growing traffic demand and subsequent traffic accidents.
@@ 1 Introduction
As reported by both the U.S. Department of Transportation (2019) and Federal Statistical Office of Switzerland (2020), motorized road traffic before the pandemic has experienced an approximate 50% growth over the past few decades. Researchers have also found that the continuous growth in traffic volume has consequently contributed to a rise in disruptive events, such as severe congestion and more frequent accidents (Chang and Xiang, 2003). As motorized individual traffic has been demonstrated with real data to recover from the pandemic (Büchel et al., 2022; Ciuffini et al., 2023), it is anticipated to persist in such increase in the coming decades (Zhang and Zhang, 2021).
@@
Meanwhile, there is a common understanding that road transportation networks can exhibit fragile properties. Fragility signifies a system’s susceptibility to exponentially escalating performance deterioration as disruptions increase in magnitude. One prominent example is the BPR function (U.S. Bureau of Public Roads, 1964), with empirical data at the link level, distinctly illustrates that travel time grows exponentially with traffic flow, leading to an infinite temporal cost when the traffic influx is at the maximal density of the network. Moreover, not limited to road transportation systems, most types of transportation systems also display similar fragile characteristics, such as railway systems (Saidi et al., 2023) and aviation systems (U.S. Congress, Office of Technology Assessment, 1984).
@@
Researchers have devoted extensive efforts to designing robust and resilient transportation systems (Shang et al., 2022; Mattsson and Jenelius, 2015). However, when accounting for the ever-growing traffic volume in urban road networks and the exponentially escalating adversarial consequences, it is natural for us to wonder whether the current level of robustness and resilience can still guarantee the performance of road networks considering a long-term time horizon. Hence, we introduce the cutting-edge concept of (anti-)fragility to explain the phenomenon that roadside performance deteriorates exponentially with linearly growing disruptions. Previous studies discussing this fragile phenomenon have primarily relied on intuitive reasoning and empirical data rather than rigorous mathematical analysis. This paper serves as a proof of concept, aiming to establish the fragile nature of road transportation systems through mathematical analysis. Additionally, as stochasticity prevails in transportation systems in the real world, we designed a numerical simulation considering real-world stochasticity to study to what extent such realistic uncertainties can influence the fragile characteristics of transportation systems. The overarching objective is to provide insights to transportation researchers for the future design of transportation systems and control strategies to be not only robust and resilient but also antifragile.
@@
The remainder of this paper is structured as follows. Section 2 introduces (anti-)fragility and reviews specific traffic-related mathematical models. Section 3 formulates the mathematical meaning of (anti-)fragility and its applications. Then we conduct the mathematical proof in Section 4, whereas Section 5 presents the numerical simulation with real-world stochasticity. With Section 6, we conclude the fragile nature and its implications for future studies.
2 Literature review
@@
Various terminologies have been proposed to evaluate the performance of transportation systems and two commonly used terms to characterize the extent of performance variations under stress are robustness (Duan and Lu, 2014) and resilience (Calvert and Snelder, 2018). However, the definitions of robustness and resilience can vary under different contexts, even within the transportation domain, due to different backgrounds and points of view of the authors (Corman et al., 2018), and sometimes even used interchangeably (Corman et al., 2018). In this study, we adopt the principle proposed in Zhou et al. (2019), wherein robustness involves evaluating a system’s ability to maintain its initial state and withstand performance degradation when confronted with uncertainties and disturbances. On the other hand, resilience emphasizes a system’s capability and promptness in recovering from major disruptions and returning to its original state. In brief, robustness relates to resistance, whereas resilience is about recovery.
@@
Nevertheless, these two terms can overlook the consideration of a longer timespan and the potential escalation of disruptions, which is particularly relevant in the context of transportation. Thus, there is the necessity to introduce a new concept to address this gap. The concept of antifragility was initially proposed by Taleb (2012) and mathematically elaborated in Taleb and Douady (2013), and it serves as a general concept aimed at transforming people’s understanding and perception of risk. By embracing current risks, we can potentially leverage and adapt to future risks with larger magnitudes. On the other hand, Ever since being proposed, antifragility has gained popularity in the risk engineering community across multiple disciplines, such as economy (Manso et al., 2020), biology (Kim et al., 2020), medicine (Axenie et al., 2022), energy (Coppitters and Contino, 2023), and robotics (Axenie and Saveriano, 2023). Generally speaking, when the performance of a system is compromised due to unexpected disruptions, the relationship between the loss in performance and the magnitude of the disruptions would be convex for a fragile system, while being concave for an antifragile system.
@@
In transportation, by unveiling the convex relationship between travel time and traffic flow with empirical data, the BPR function (U.S. Bureau of Public Roads, 1964), as shown in Eq. 1 gives an intuitive example showing the fragility of the transportation systems on a link level. However, to uphold the statement that road transportation systems are fragile, an empirical function like BPR alone is not sufficient without rigid mathematical proof. Also, it is desired to show the fragility not only on the link level but also on the macroscopic level, as well as for different types of disruption.
@@
Prior to diving into the demonstration of the fragile characteristics of road transportation systems, various traffic-related models should first be explained. The assessment of traffic performance can be conducted at either the microscopic or macroscopic level. However, due to the non-identical properties of traffic across these levels, researchers have formulated diverse traffic models to offer a more precise description of traffic dynamics at different levels. These models serve the purpose of accurately portraying traffic behavior, with some being derived numerically from data, while others are developed through analytical methods.
Although various traffic models on the microscopic level have been proposed, we select two most representative models, one numerical and one analytical, as examples. The earliest study on traffic performance was carried out and described in Greenshields et al. (1934) on a section of a highway and created the first Fundamental Diagram (FD), as shown in Fig. 1(a), which exhibits the relationship between traffic flow, denoted as $q$, and density, denoted as $k$, in the shape of a second-degree polynomial. Later, other researchers also developed FDs in various forms, with the most commonly applied FD being proposed in Daganzo (1994), as shown in Fig. 1(b), and is characterized by two linear functions. While the Greenshields FD is entirely generated through the fitting of the polynomial coefficients, the Daganzo FD, on the other hand, is derived analytically and incorporates variables with physical meanings, i.e., the free-flow speed, back-propagation speed, and critical density, denoted as $v_{f}$, $w$, and $k_{c}$, respectively.
@@
On the macroscopic level, a Macroscopic Fundamental Diagram (MFD) can be produced by aggregating data points gathered from each single link within this network. Similar to FD, MFD can be represented through different models, and the most representative and widely applied approaches include polynomials and multi-regime linear functions. Daganzo and Geroliminis (2008) is the first study to ever analytically generate an MFD by using the Method of Cuts (MoC). This analytical method demonstrated the feasibility of using variables with physical meanings to approximate an MFD with multi-regime linear functions in a homogenous region, as shown in Fig. 2(b). Leclercq and Geroliminis (2013) modified the original MoC to accommodate topology and signal timing heterogeneity within the network. Although theoretically an infinite number of cuts can be generated to approximate the MFD, only the practical cuts, the solid lines in Fig. 2(b), are used for simplicity reasons.
@@
As also acknowledged in these works, some variables, particularly the parameters related to signalization, are hard to acquire in the real world with actuated signals. Therefore, simplified MFDs are also widely adopted when solving more applicational problems like traffic control. Daganzo et al. (2018) approximated a simplified MoC-based MFD using only three cuts from one of each stationary, forward, and backward observer, and formed the uMFD. Other researchers also approximated the MFDs based on collected data with polynomials, e.g., a third-degree polynomial in Geroliminis et al. (2013) as shown in Fig. 2(a). It should be noted that since the data points near the maximal density are scarce overall, the approximated MFDs are sometimes not strictly realistic. For example, the third-degree polynomial in Geroliminis et al. (2013) has only one real root, indicating the speed is above zero despite that the network has reached its maximal density and formed a gridlock. As claimed in Daganzo and Geroliminis (2008), since MFD should be concave, the flow and speed are both zero by definition at maximal density.
@@ 3 Problem formulation
As explained in Section 2, an (anti-)fragile response of a system can be characterized through a nonlinear relationship between the performance loss and the magnitude of the disruption, as shown in Fig. 3(a). Both nonlinear functions can be represented by Jensen’s inequality (Jensen, 1906), with either $E[g(X)]\geq g(E[X])$ for a fragile response or $E[g(X)]\leq g(E[X])$ for an antifragile response. This relationship can then be determined through the second derivative (Ruel et al., 1999), i.e., a positive second derivative featuring a convex function and hence a fragile system. It should be noted that the calculation of the derivatives is only possible when the function is continuous and differentiable, which means the underlying mathematical model should be present.
However, in most real-world scenarios, the mathematical function of the system is unknown, and only discrete measurements of the system’s performance are available. In this case, we can calculate the distribution skewness to determine the (anti-)fragile property of the system, as illustrated with an example of financial deficit in Taleb and Douady (2013) and Coppitters and Contino (2023) when designing an antifragile renewable energy system. A negative skewness represents the long tail points left and indicates an antifragile response, as shown in Figure 3(b).
@@
In this paper, we address three sets of opposing concepts when studying the fragile nature of road transportation systems:
@@
As microscopic and macroscopic traffic models were introduced in Section 2, in this section, we initiate the discussion by differentiating between demand and supply disruptions, with the latter also sometimes referred to as MFD disruptions. As traffic networks primarily involve the management of supply and demand, we consider that any traffic disruption in the real world can be classified as either a demand or a supply disruption. A demand disruption can be easily understood as, for example, surging traffic due to a social event, whereas a supply disruption may indicate an impaired network performance due to external factors, such as adversarial weather or lane closure. Additionally, since disruptions represent abnormal cases that only exist temporarily, we consider not only the onset of disruptions but also assess the recovery process of the systems following such disruptions. The scheme of onset and recovery from either demand or supply disruptions are illustrated in Figure 4(a) and Figure. 4(b) on the macroscopic level. Since FDs share similar profiles as MFDs, we use MFDs here as examples for illustration. We denote the FD/MFD profile as $G(k)$ and assume a constant base demand in the network as $q$. The initial density corresponding to the base demand, critical density, the new density after disruption, and the gridlock density are donated as $k_{0}$, $k_{c}$, $k^{\prime}$, and $k_{\rm max}$, respectively. For supply disruption, we introduce a disruption magnitude coefficient $r$ so that the disrupted MFD profile can be represented as $(1-r)G(k)$. It should also be noted that on the network level, instead of traffic flow - density, MFD can also be represented with vehicle accumulation - trip completion, such as in Zhou and Gayah (2021); Kouvelas et al. (2017); Genser and Kouvelas (2022). Hence, with vehicle accumulation denoted as $n$, $G(k)$ and $k_{*}$ can also be replace with $G(n)$ and $n_{*}$.
@@
As the study of (anti-)fragility also involves the recovery process following a potential disruption, several assumptions need to be established to define the scope of our study. A critical condition to avoid is the network succumbing to a complete gridlock, where recovery is not possible anymore.
  1. 1.
    For demand disruptions, $k^{\prime}>k_{c}$, whereas for supply disruptions, $k^{\prime}<k_{c}$.
    A surging demand is considered a disruption only when its presence leads to a reduction in the network’s serviceability, i.e., causing the traffic state to enter the congested zone of the MFD. On the other hand, for supply disruption with the constant base demand, if the traffic state can ever surpass the maximal capacity of the disrupted MFD profile, then the density will continue to accumulate until the network reaches gridlock, as the incoming flow is greater than the maximal possible outgoing flow.
  2. 2.
    For demand disruptions, $q<G(k^{\prime})$, whereas for supply disruptions, $q<(1-r)G(k_{c})$.
    Likewise, the necessity of this assumption also lies in the avoidance of gridlock for both demand and supply disruptions. If the base demand is higher than the outgoing flow after a demand disruption or the maximal flow after a supply disruption, then the traffic state will continue to move to complete gridlock.
  3. 3.
    The onset of the disruptions can happen instantaneously, while the recovery from disruptions is a gradual process.
@@
    This can be particularly true for demand disruptions characterized by a rapid influx of traffic into the network. Therefore, it can be understood as a sudden change from one stable traffic state to another stable state. On the other hand, the recovery from disruptions can take significantly longer and thus it is a traffic unloading and congestion dissipation process instead of an instantaneous event.
4 Mathematical proof of the fragility of road transportation systems

### 04 Mar 2024 v2

@@ Abstract
Major cities worldwide experience problems with the performance of their road transportation systems, and the continuous increase in traffic demand presents a substantial challenge to the optimal operation of urban road networks and the efficiency of traffic control strategies. Although robust and resilient transportation systems have been extensively researched over the past decades, their performance under an ever-growing traffic demand can still be questionable. The operation of transportation systems is widely believed to display fragile property, i.e., the loss in performance increases exponentially with the linearly increasing magnitude of disruptions, which undermines their continuous operation. Nowadays, the risk engineering community is embracing the novel concept of antifragility, which enables systems to learn from historical disruptions and exhibit improved performance as disruption levels reach unprecedented magnitudes. In this study, we demonstrate the fragile nature of road transportation systems when faced with demand or supply disruptions. First, we conducted a rigorous mathematical analysis to establish the fragile nature of the systems theoretically. Subsequently, by taking into account real-world stochasticity, we implemented a numerical simulation with realistic network data to bridge the gap between the theoretical proof and the real-world operations, to reflect the potential impact of uncertainty on the fragile property of the systems. This work aims to demonstrate the fragility of road transportation systems and help researchers better comprehend the necessity to explicitly consider antifragile design for future traffic control strategies, coping with the constantly growing traffic demand and subsequent traffic accidents.
@@ 1 Introduction
As reported by both the U.S. Department of Transportation (2019) and Federal Statistical Office of Switzerland (2020), motorized road traffic before the pandemic has experienced an approximate 50% growth over the past few decades. Researchers have also found that the continuous growth in traffic volume has consequently contributed to a rise in disruptive events, such as severe congestion and more frequent accidents (Dickerson et al., 2000; Chang and Xiang, 2003). With evidence confirming the recovery of individual motorized traffic from the pandemic (Büchel et al., 2022; Marra et al., 2022; Ciuffini et al., 2023), it is expected that this upward trend will continue in the coming decades (Zhang and Zhang, 2021).
@@
Meanwhile, there is a common understanding that road transportation networks can exhibit fragile properties. Fragility signifies a system’s susceptibility to exponentially escalating performance deterioration as disruptions increase in magnitude. One prominent example of such fragile characteristics is the BPR function (U.S. Bureau of Public Roads, 1964), which distinctly illustrates with empirical data at the link level that travel time grows exponentially with traffic flow, leading to an infinite temporal cost when the traffic influx is at the maximal density of the network. Moreover, not limited to urban road networks, many other types of transportation systems can also display similar fragile responses to increasing disruption levels, such as in railway (Gibson et al., 2002; Saidi et al., 2023) and aviation systems (U.S. Congress, Office of Technology Assessment, 1984).
@@
Researchers have devoted extensive efforts to the assessment and design of robust and resilient transportation systems of all kinds, such as in railway systems (Corman et al., 2014; Larsen et al., 2014), public transportation operation (Cats, 2016; Fuchs and Corman, 2019), aviation (Isaacson et al., ), and road networks (Ampountolas et al., 2017; Yang et al., 2019; Leclercq et al., 2021). However, when accounting for the ever-growing traffic volume in urban road networks and the exponentially escalating adversarial consequences, it is natural for us to wonder whether the current level of robustness and resilience can still guarantee the performance of road networks considering a long-term time horizon. Hence, we introduce the cutting-edge concept of antifragility to explain the phenomenon that roadside performance deteriorates exponentially with linearly growing disruptions. Previous studies discussing the fragile response, such as the BPR function, have primarily relied on empirical data and intuitive reasoning rather than rigorous mathematical proof. This paper, on the other hand, serves as a proof of concept, aiming to establish the fragile nature of road transportation systems through mathematical analysis. Additionally, as stochasticity prevails in transportation systems in the real world, we also designed a numerical simulation considering real-world stochasticity, to study to what extent such realistic uncertainties can influence the fragile characteristics of transportation systems. The overarching objective of this paper is to provide insights to transportation researchers for the future design of transportation systems and control strategies to be not only robust and resilient but also antifragile.
@@
The remainder of this paper is structured as follows. Section 2 introduces antifragility and reviews specific traffic-related mathematical models. Section 3 formulates the mathematical definition of (anti-)fragility and its applications. Then we conduct the mathematical proof in Section 4, whereas Section 5 presents the numerical simulation with real-world network and stochasticity. With Section 6, we conclude the fragile nature of road transportation systems and its implications for future studies.
2 Relevant Literature
@@
Various terminologies have been proposed to evaluate the performance of road transportation systems, including their control strategies, and two commonly used terms to characterize the extent of performance variations under stress are robustness (Duan and Lu, 2014; Shang et al., 2022) and resilience (Mattsson and Jenelius, 2015; Calvert and Snelder, 2018). However, the definitions of robustness and resilience can vary under different contexts, even within the transportation domain itself, and are sometimes used interchangeably due to the different backgrounds and points of view of the authors (Tang et al., 2020; Corman et al., 2018). In this study, we adopt the definition proposed in Zhou et al. (2019), wherein robustness involves evaluating a system’s ability to maintain its initial state and withstand performance degradation when confronted with uncertainties and disturbances. On the other side, resilience emphasizes a system’s capability and promptness in recovering from major disruptions and returning to its original state. In brief, robustness relates to resistance, whereas resilience is about recovery.
@@
Nevertheless, these two terms can overlook the consideration of a longer timespan and the potential escalation of disruptions, which is particularly relevant in transportation as the traffic demand grows continuously and consequently so as the occurrences of accidents. Thus, there is the necessity to introduce a new term to address this gap. The novel concept of antifragility was initially proposed in Taleb (2012) and mathematically elaborated in Taleb and Douady (2013); Taleb and West (2023), and it serves as a general concept aimed at transforming people’s understanding and perception of risk. By embracing current risks, we can potentially leverage and adapt to future risks of greater magnitudes. When employed in systems and control, (anti-)fragility can be conceptualized as a nonlinear relationship between the performance and the magnitude of disruptions. If the performance is compromised due to unexpected disruptions, the relationship between the loss in performance and the disruptions would be convex for a fragile system, while being concave for an antifragile system. Ever since being proposed, antifragility has gained popularity in the risk engineering community across multiple disciplines, such as economy (Manso et al., 2020), biology (Kim et al., 2020), medicine (Axenie et al., 2022), energy (Coppitters and Contino, 2023), robotics (Axenie and Saveriano, 2023), and lately in transportation (Sun et al., 2024). It should also be highlighted that although systems can be fragile by nature, proper intervention and control strategies can enhance their antifragility against increasing disruptions (Axenie et al., 2023).
@@
In transportation, by unveiling the convex relationship between travel time and traffic flow with empirical data, the BPR function (U.S. Bureau of Public Roads, 1964), as shown in Eq. 1 has given an intuitive example showing the fragility of the transportation systems on a link level. Together with its variations (Dowling and Skabardonis, 1993; Skabardonis and Dowling, 1997), the BPR function has been extensively applied in the estimation of the link (route) travel time (Lo et al., 2006; Ng and Waller, 2010; Wang et al., 2014). However, to uphold the statement that road transportation systems are fragile in general, an empirical function like BPR alone is not sufficient without rigid mathematical proof. It is also desired to show the fragility more broadly, i.e., not only on the link level but also on the macroscopic level, as well as for different types of disruptions.
@@
The assessment of traffic performance can be conducted at either the microscopic or macroscopic level, sometimes also mesoscopically (Zehe et al., 2015). However, due to the non-identical traffic characteristics across different levels, researchers have formulated diverse models to offer a more precise description of traffic dynamics on different levels. While some models are generated numerically from data, others are derived through analytical methods (Mariotte et al., 2017).
The earliest study on traffic performance was carried out and described in Greenshields et al. (1934) on a section of a highway and yielded the first Fundamental Diagram (FD), as shown in Fig. 1(a), which exhibits the relationship between traffic flow, denoted as $q$, and density, denoted as $k$, in the shape of a second-degree polynomial. Later, other researchers also developed FDs in various forms, with one of the most commonly applied FDs being proposed in Daganzo (1994), as shown in Fig. 1(b), which is characterized by two linear functions. While the Greenshields FD is entirely generated through fitting the polynomial coefficients with on-site data, the Daganzo FD, on the other hand, is derived analytically by treating the traffic flow hydrodynamically and incorporates variables with physical meanings, i.e., the free-flow speed, back-propagation speed, and critical density, denoted as $v_{f}$, $w$, and $k_{c}$, respectively.
@@
On the network level, with the assumption of a homogeneous region, a Macroscopic Fundamental Diagram (MFD) can be produced by aggregating data points gathered from representative links within this network. Similar to FDs, MFDs can be approximated through different models, and the most widely applied approaches include polynomials and multi-regime linear functions. MFDs can be fit numerically from field measurements as a cubic polynomial, as in Haddad and Shraiber (2014); Sirmatel and Geroliminis (2018) and illustrated in Fig. 2(a), and therefore, it’s commonly seen in research works simulating traffic control on a macroscopic level with a given MFD. On the other hand, based on the framework of variation theory (Daganzo, 2005), Daganzo and Geroliminis (2008) is the first study to analytically generate an MFD, and is often referred to as the Method of Cuts (MoC). Instead of installing loop detectors under the pavement and gathering massive traffic flow data, MoC can be applied to derive an MFD directly from roadside variables with physical meanings, such as free flow speed, traffic signal cycle, lane length, etc. An example of multi-regime linear functions MFD from MoC is shown in Fig. 2(b). Leclercq and Geroliminis (2013) modified the original MoC to accommodate topology and signal timing heterogeneity within the network. Although theoretically an infinite number of cuts can be generated to approximate the MFD, only the practical cuts, the solid lines as shown in Fig. 2(b), are used for simplicity reasons. Although some other analytical methods have also been proposed to produce an MFD, such as through stochastic approximation (Laval and Castrillón, 2015), Tilg et al. (2020) has demonstrated that MoC yields a more accurate upper bound for the MFD. Lately, researchers have also explored the potential utilization of MFD in railway (Corman et al., 2019) and aviation operations (Safadi et al., 2023), which yields some positive possibilities of extending the MFD into other modes.
@@
As also acknowledged in these works, some variables, particularly the parameters related to signalization, can be hard to acquire in the real world with actuated signals. Therefore, simplified MFDs are also widely adopted when solving more applicational problems in the real world. Daganzo et al. (2018) approximated a simplified MoC-based MFD using only three cuts from one of each stationary, forward, and backward observer, and formed the uMFD. Another shortcoming of the numerical-based MFD is that, since the data points near the maximal density are scarce overall, the approximated MFD can lose its fidelity near the maximal density. For example, the third-degree polynomial in Geroliminis et al. (2013) has only one real root, indicating the speed is above zero despite the fact that the network has already reached a gridlock, which cannot be realistic. As claimed in Daganzo and Geroliminis (2008), since MFDs should be concave, the flow and speed are both zero by definition at maximal density.
@@ 3 Problem formulation
As explained in Section 2, an (anti-)fragile response of a system can be characterized through a nonlinear relationship between the performance loss and the magnitude of the disruption, as shown in Fig. 3(a). Both nonlinear functions can be represented by Jensen’s inequality (Jensen, 1906), with either $E[g(X)]\geq g(E[X])$ for a fragile response or $E[g(X)]\leq g(E[X])$ for an antifragile response. This relationship can then be determined through the second derivative (Ruel et al., 1999), i.e., a positive second derivative featuring a convex function and hence a fragile system and vice versa. It should be noted that the calculation of the derivatives is only possible when the function is continuous and differentiable, which means the underlying mathematical model representing the system needs to be known beforehand.
However, in most real-world scenarios, the mathematical function of the system is agnostic, and only discrete measurements of the system’s performance are available. In this case, we can calculate the distribution skewness to determine the (anti-)fragile property of the system, as illustrated with an example of financial deficit in Taleb and Douady (2013) and Coppitters and Contino (2023) when designing an antifragile renewable energy system. A negative skewness represents the long tail pointing to the left and indicates an antifragile response, as shown in Fig. 3(b).
@@
In this paper, we address three sets of opposing concepts for the mathematical analysis when studying the fragile nature of road transportation systems:
@@
As microscopic and macroscopic traffic models were introduced in the previous Section 2, here we start with differentiating between demand and supply disruptions, with the latter sometimes referred to as MFD disruptions. Since traffic networks primarily involve the management of supply and demand, we consider that any traffic disruption in the real world can be classified as either a demand or a supply disruption. A demand disruption can be easily understood as, for example, surging traffic due to a social event, whereas a supply disruption may indicate an impaired network due to external factors, such as adversarial weather or lane closure. Additionally, since disruptions represent abnormal cases that only exist temporarily, we consider not only the onset of disruptions but also assess the recovery process of the systems following such disruptions. This coincides with the aforementioned definitions of robustness and resilience, i.e., robustness is about resistance against disruptions while resilience is about the recovery from them. By considering both the onset and recovery processes, it becomes possible to compare the performance between either robust and antifragile designs or resilient and antifragile designs. The scheme of onset and recovery from demand or supply disruptions are illustrated in Fig. 4(a) and Fig. 4(b), and since FDs share similar profiles as MFDs, we simply use MFDs here as examples for illustration. We denote the FD/MFD profile as $G(k)$ and assume a constant base demand in the network as $q$, resulting in an equilibrium traffic state in the network in the absence of any disruption. The initial density at equilibrium, the critical density, the new density after disruption, and the gridlock density are denoted as $k_{0}$, $k_{c}$, $k^{\prime}$, and $k_{\rm max}$, respectively. For the study of supply disruptions, we introduce a disruption magnitude coefficient, denoted as $r$, so that the disrupted MFD profile can be represented as $(1-r)G(k)$. It should also be noted that on the network level, instead of traffic flow - density, MFD can also be represented with vehicle accumulation - trip completion, such as in Zhou and Gayah (2021); Kouvelas et al. (2017); Genser and Kouvelas (2022). Hence, with vehicle accumulation denoted as $n$, $k_{*}$ and $G(k)$ can also be replace with $n_{*}$ and trip completion $G(n)$.
@@
Several assumptions need to be established to define the scope of our work. Also, as the study of fragility involves the recovery process from disruptions, a critical condition to avoid here is the network succumbing to a complete gridlock, where recovery is not possible anymore.
Assumption 1.
We study the onset of disruptions focusing only on the two stable traffic states before and after the disruptions, while the recovery is regarded as a gradual process of congestion dissipation.
@@
For the onset of demand disruptions, we assume a rapid influx of traffic into the network, which can be characterized as an instantaneous event. For the onset of supply disruptions, although it may take some time for the number of vehicles within the network to accumulate, we understand it as a change from one stable traffic state $G(k_{0})$ to the final stable state $(1-r)G(k^{\prime})$, i.e., the equilibrium points in blue as shown in Fig. 4(b). The process of recovery from demand or supply disruptions, on the other hand, takes significantly longer as it is a traffic unloading and congestion dissipation process instead of an instantaneous event.
Assumption 2.
For demand disruptions, we assume $k^{\prime}>k_{c}$, whereas $k^{\prime}<k_{c}$ for supply disruptions.
A surging demand should be considered a disruption only when its presence leads to a reduction in the network’s maximal possible serviceability, i.e., causing the traffic state to enter the congested zone of the MFD and the flow is below the network capacity. On the other hand, for supply disruption with the constant base demand, if the traffic state can ever surpass the maximal capacity of the disrupted MFD profile, indicating the incoming flow is even greater than the maximal capacity of the network after supply disruptions, then the traffic density will continue to accumulate until the network reaches a full gridlock, and there will be no equilibrium point in this case after a disruption, violating Assumption 1.
Assumption 3.
For demand disruptions, we assume $q<G(k^{\prime})$, whereas $q<(1-r)G(k_{c})$ for supply disruptions.
Likewise, the necessity of this assumption also lies in the avoidance of gridlock for both demand and supply disruptions. If the base demand is higher than the outgoing flow after a demand disruption or the maximal capacity of the network after a supply disruption, then the traffic state will continue to move to complete gridlock.
4 Mathematical analysis of the fragility of road transportation systems